# Topological flow data analysis for transient flow patterns: a graph-based approach

### v1

@@ 3.1.1. Transition of streamline topology
Figure 9. Schematic of topological transitions of local flow patterns at the top left corner through marginal structurally unstable patterns. (a) The flow pattern at $t=t_{140}$ represented by $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$. (b) A structurally unstable flow pattern with heteroclinic connections. (c) The flow pattern at $t=t_{300}$ represented by $c_{+}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$. (d) A structurally stable flow pattern represented by $c_{+}(\sigma_{+},c_{-}(\sigma_{-}))$, which is expected to exist between the patterns at $t=t_{300}$ and $t=t_{360}$ according to Sakajo and Yokoyama (2015). (e) A structurally unstable flow pattern with a degenerate boundary saddle. (f) The structurally stable pattern at $t=t_{360}$ represented by $c_{+}(\sigma_{+})$.
3.1.2. Transition diagram of the flow evolution
As we see in Section 3.1.1, the change of flow patterns from $t=t_{0}$ to $t=t_{580}$ at $Re=14000$ is expressed as a cycle of the COT representations of the five topologically equivalent patterns in Table 2. Those COT representations share the same root structure $\beta_{\emptyset-}$ followed by $\sigma_{-}$ expressing the large clockwise rotational region in the cavity. Hence, the five topological equivalence classes are distinguished by the sequences of the COT symbols $c_{+}^{i}$ ($i=0,1,2$) to the right of $\sigma_{-}$, which represents the topological structure in the recirculating flow domains at the cavity corners.
We now track the change of the sequences of $c_{+}^{i}$ in the COT representations in the temporal direction as follows. Applying TFDA to the flow pattern at each time, we extract the sequence of $c_{+}^{i}$ in the COT representation. We do nothing as long as this sequence remains the same. When the sequence changes, we create a directed edge connecting the nodes labelled by the COT sequences before and after the change. Repeating the procedure for the whole evolution, we obtain a directed graph representing the transition between the nodes labelled by the COT sequences. We call this directed graph a transition diagram. By creating the transition diagram, we reduce the evolution of the lid-driven cavity flow into a discrete dynamical system among the finite number of topologically equivalent flow patterns. Figure 10(a) is the transition diagram constructed by the flow evolution at $Re=14000$ from $t=t_{0}$ to $t_{10000}$. The label of each node in this diagram represents a sequence of COT symbols. For instance, the label “c+0(b++,c-).c+1.c+1.c+2” corresponds to the sequence $c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})$ in the COT representation at $t=t_{360}$ in Table 2.
In addition, from the transition diagram, we obtain an $N\times N$ transition matrix $\mathcal{M}$ between $N$ nodes as follows. For each node $i$, we count the number of transitions to the node $j(\neq i)$, say $C_{ij}$, for $j=1,\dots,N$. Note that $C_{ii}=0$ since we do not count the transition from the node $i$ to itself when constructing the transition diagram. The entries of $\mathcal{M}$ are then given by
${\mathcal{M}}_{ij}=\frac{C_{ij}}{\langle C\rangle_{i}},\qquad\langle C\rangle_{i}=\sum_{j=1}^{N}C_{ij}.$
Then we have $\sum_{j=1}^{N}{\mathcal{M}}_{ij}=1$ for any $i$. The eigenvalues of this transition matrix $\mathcal{M}$ provide information about the discrete dynamical system as the Markov process on the graph. Let us consider the initial probability distribution $\mu(0)$ on the graph. Suppose that the transition matrix $T$ has the distinct eigenvalues $\lambda_{j}$ and its associated left eigenvector ${\bm{\ell}}_{j}$ for $j=1,\dots,N$. Recall that all eigenvalues of the transition matrix exist in the unit disk. Then the probability distribution at time step $k$ is represented by
(9) $\mu(k)^{T}=\sum_{j=1}^{m}d_{j}\lambda_{j}^{k}{\bm{\ell}}_{j}^{T}+\sum_{j=m+1}^{N}d_{j}\lambda_{j}^{k}{\bm{\ell}}_{j}^{T}$
with the constants $d_{j}$ satisfying $\mu(0)^{T}=\sum_{j=1}^{N}d_{j}{\bm{\ell}}_{j}^{T}$, where $|\lambda_{j}|=1$ for $j=1,\dots,m$ and $|\lambda_{j}|<1$ for $j=m+1,\dots,N$. The first term represents the persistent dynamics for all time steps, while the second term is the transient part decaying as time proceeds. Figure 10(b) shows the distribution of the eigenvalues of the transition matrix for $Re=14000$. Recall that the transition matrix always has the Perron–Frobenius eigenvalue $1$. The other four eigenvalues are located at the vertices of the right pentagon on the unit circle. Hence, the decomposition (9) indicates that this discrete dynamical system is in periodic motion with period five.
Figure 10. (a) Transition diagram for the evolution of the lid-driven cavity flow at $Re=14000$. (b) Plot of the eigenvalues of the transition matrix in the complex plane.
3.2. Comparison across Reynolds numbers
3.2.1. Graph analysis of transition diagram
We investigate the behaviour of the reduced discrete dynamical system of the lid-driven cavity flow in detail using techniques of discrete graph analysis stated above. Let us consider the range of Reynolds numbers from $Re=14000$ to $Re=16000$, where periodic behaviour changes to chaotic aperiodic behaviour according to Shen (1991).
Figures 11(a-c) are the transition diagrams at $Re=14500$, $Re=15000$, and $Re=15500$, respectively. Figure 12 shows the distributions of the eigenvalues of the transition matrix corresponding to the transition diagrams for various Reynolds numbers. At $Re=14500$, the diagram forms a hexagonal transition among six topologically equivalent flow patterns in one direction. The eigenvalues of the corresponding transition matrix in Figure 12(b) are located at the vertices of the right hexagon on the unit circle, indicating that the motion of the discrete dynamical system is periodic with period six. When the Reynolds number becomes $Re=15000$, the transition diagram in Figure 11(b) consists of two triangles. It suggests that the transitions of the local flow patterns at the recirculating regions $c_{+}^{0}$ and $c_{+}^{1}$ occur alternately. Figure 12(d) shows that the eigenvalues except for the Perron–Frobenius eigenvalue are inside the unit circle. Since the arguments of the two eigenvalues are ±$\pm\frac{2}{3}\pi$, a decaying periodic motion with period three exists. The remaining one is near the origin, indicating the degenerate node “c0.c1.c2” connects the two triangles in the transition diagram. When the Reynolds number gets larger, $Re=15500$, the transition diagram becomes complicated as shown in Figure 11(c). Then we find no eigenvalues generating almost periodic behaviour as shown in Figure 12(f). Hence, we expect the change from periodic motion to aperiodic one between $Re=15000$ and $Re=15500$.
Figure 11. Transition diagrams for the lid-driven cavity flow at (a) $Re=14500$, (b) $Re=15000$, and (c) $Re=15500$.
Figure 12. Eigenvalues of the transition matrices corresponding to the lid-driven cavity flow in the transition regime of the Reynolds number. (a) $Re=14250$. (b) $Re=14500$. (c) $Re=14750$. (d) $Re=15000$. (e) $Re=15250$. (f) $Re=15500$. (g) $Re=15750$. (h) $Re=16000$.
[…]

### v2

@@ 3.1. Transition of streamline topology
Figure 10. Schematic of topological transitions for $Re=14000$ of local flow patterns in the top left corner from $t=t_{300}$ to $t=t_{400}$ that is interpolated based on the topological transition theory Sakajo and Yokoyama (2015). (a) The flow pattern at $t=t_{300}$ represented by $c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})$. (b) A singular orbit with a degenerate pinching emerges. (c) Creation of a local self-connected saddle connection between the two boundary saddle connections through the pinching. This is structurally stable, but it is not detected by TFDA from the present dataset. (d) A structurally unstable flow pattern with heteroclinic connections. (e) The structurally stable flow pattern observed at $t=t_{360}$ represented by $c_{+}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))$. (f) The inner clockwise boundary saddle connection $c_{-}(\sigma_{-})$ disappears through the pinching transition on the left wall as shown in Figure 9. (g) A structurally stable represented by $c_{+}(b_{++}\{\sigma_{+},\sigma_{+}\})$ that is not detected in the dataset by TFDA, either. (h) The local orbit structure $b_{++}$ disappears through a pinching transition. (i) The structurally stable pattern represented by $c_{+}(\sigma_{+})$ at $t=t_{400}$.
It is important to note that a substantial body of literature has been devoted to the study of topological bifurcations of this kind from analytical perspectives; see, e.g., Brøns et al. (1999); Brøns (2007); Bakker (1991); Brøns et al. (2007). These studies consider certain perturbation parameters that develop rigorous local bifurcations (e.g., bubble (or zone) creations and global bifurcations) within parameter spaces. On the other hand, the pattern change presented here is one possible route through the minimum number of transitions suggested by topological transition theory. Hence, to determine what transitions actually occur, it is necessary to perform TFDA on time series data sampled with a finer temporal frequency. Our topological approach, along with that of Fomenko (2004) in the field of integrable systems, focuses on global topological structures, with the aim of providing a comprehensive classification of all possible transitions that can occur combinatorially.
3.2. Transition diagram of the flow evolution
As we see in Section 3.1, the change of flow patterns from $t=t_{0}$ to $t=t_{580}$ is expressed as a COT representation cycle. Topologically equivalent flow patterns are represented by a single COT representation, which we refer to as a topological state. Under this representation, the continuous evolution of the flow can be reduced to a discrete dynamical process on the set of topological states. To construct this description, we compute the COT representation of the streamfunction at each time step during the evolution. This produces a time series of COT representations that captures the sequence of topological configurations realised by the flow. Each distinct COT representation is identified as a node corresponding to a topological state. Whenever the representation changes from one time step to the next, a directed edge is introduced from the preceding node to the subsequent one. Through this procedure, we obtain a directed graph whose nodes represent topological states and whose edges represent transitions between them. We refer to this graph as a transition diagram. It provides a compact visual representation of the evolution of the lid-driven cavity flow in terms of discrete topological changes.
Figure 11 presents the transition diagram for $Re=14000$, constructed by tracking the flow evolution over an extended time interval from $t=t_{0}$ to $t_{10000}$. Each node is labelled by a simplified COT expression. For example, the label $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(b_{++},c_{-})\cdot c_{+}^{1}\cdot c_{+}^{1}\cdot c_{+}^{2}\})$ corresponds to the full COT representation $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$. We note that the topological states corresponding to these nodes are topologically equivalent to the patterns in panels (a)–(f) of Figure 8. The size of each node in Figure 11 indicates the time spent on the corresponding state, while the thickness of each edge represents the number of observed transitions between the connected states. The diagram shows that the topological state $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}\cdot c_{+}^{1}\cdot c_{+}^{2}\})$, which is the simplest topological structure as in Figure 8(a), has the longest existence time in the time series. As observed in Table 2, there is a dominant hexagonal one-way transition cycle with thick directed edges, corresponding to the transition of topological states in panels (a) $\to$ (b) $\to$ (c) $\to$ (d) $\to$ (e) $\to$ (f) $\to$ (a) of Figure 8. This constitutes the principal transition route. In addition, a short-cut transition is visible, connecting $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}\cdot c_{+}^{0}\cdot c_{+}^{1}\cdot c_{+}^{2}\})$ directly to $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(b_{++},c_{-})\cdot c_{+}^{1}\cdot c_{+}^{1}\cdot c_{+}^{2}\})$. This arises because the intermediate state $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}\cdot c_{+}^{0}\cdot c_{+}^{1}\cdot c_{+}^{1}\cdot c_{+}^{2}\})$ appears for at most one frame at the present sampling rate and is occasionally missed, as indicated in Table 2. It should also be noted that two instances of $c_{+}^{1}$ are visually identifiable in Figure 8(b), yet they fall below the adopted detection threshold, whereas in Figure 8(c) they are counted as two distinct $c_{+}^{1}$ components. Therefore, we emphasize that the structure of the transition diagram depends on the threshold and the sampling frequency.
Figure 11. Transition diagram for the evolution of the lid-driven cavity flow for $Re=14000$.
4. Topological flow data analysis across Reynolds numbers
We demonstrate how TFDA extracts physical information from the lid-driven cavity flow by exploiting the time series of COT representations. Our analysis focuses on Reynolds numbers in the range of $14000\leq Re\leq 16000$, where the flow is considered to have reached a statistically stationary state at time $t=T_{s}=5000$. From this time onward, we analyse the subsequent evolution of the flow topology. Specifically, we sample the time series of the streamfunction over the interval $0<t-T_{s}\leq 200$. Snapshots are recorded every $10\times\Delta t=0.01$, where $\Delta t=10^{-3}$ is the time-step size used in the direct numerical simulation. This yields a total of $20000$ streamfunction snapshots at $t_{k}=T_{s}+0.01\times k$ for $k=1,\ldots,20000$ for each time series. The Reynolds numbers are sampled non-uniformly within the interval: every $250$ for $14000\leq Re\leq 15250$, and every $50$ for $15300\leq Re\leq 16000$, where the finer resolution in the higher range is used to capture more detailed dynamical changes.
4.1. Transition of flow complexity with topological states
We investigate the variation of topological states as a function of the Reynolds number. In particular, we show that the increase in the number of topological states, or nodes in the transition diagram, has a critical behaviour around the transition from the periodic to the quasi-periodic state. We first visualise transition diagrams in the range of $15000\leq Re\leq 16000$, where the periodic behaviour changes to chaotic aperiodic behaviour according to Shen (1991). Figure 12(a) presents the transition diagram for $Re=15000$. The nodes are not labelled by COT representations but ID numbers instead due to insufficient display space. The correspondence between ID numbers and COT representations is listed in Table 4. These IDs function simply as numeric identifiers of topological states here and will be used consistently in subsequent sections. The diagram shows that the flow is periodic and the evolution follows the sequence $1\to 2\to 8\to 9\to 1\to 5\to 10\to 11\to 10\to 12\to 13\to 1$. This constitutes a cycle of length $11$. However, since certain states share the same COT representation, the diagram appears to contain three distinct cycles. When the Reynolds number is $Re=15500$, the transition diagram becomes intricate, as shown in Figure 12(b). It indicates that the flow changes from periodic to aperiodic behaviour between $Re=15000$ and $15500$. Figure 12(c) presents the transition diagram for $Re=16000$, where a larger number of topological states are observed.
To quantify these regime changes in dynamics complexity, we simply count the number of distinct topological states $n_{top}(Re)$, as a function of the Reynolds number. The resulting dependence is shown in Figure 13. The quantity $n_{top}$ shows a sharp increase just below $Re=15500$, suggesting critical behaviour and serving as an order parameter for the change in the flow regime. We fit this rapid growth with a power-law function, as illustrated in the inset of Figure 13. With a critical Reynolds number $Re_{c}=15400$, we obtain a reasonable fit of the form $n_{top}(Re)\sim(Re-Re_{c})^{0.10}$. It may also be of interest to determine the asymptotic scaling of $n_{top}(Re)$ as $Re\to\infty$, but this is beyond the scope of this article. Since a COT has a tree structure, as illustrated in Figure 7, the number of possible representations can grow exponentially in this regime.
Figure 12. Transition diagrams for the lid-driven cavity flow for (a) $Re=15000$, (b) $Re=15500$, and (c) $Re=16000$.
It should be noted that $n_{top}(Re)$ depends on the threshold $\varepsilon$ of TFDA, below which the spatial variations of the streamfunction are neglected, as described in Section 3.1. Hence, quantitative aspects of critical behaviour, such as the exponent value, also depend on the threshold $\varepsilon$. Nevertheless, we consider that the dependence becomes weak if the threshold is taken to be sufficiently small for a given Reynolds number.
Figure 13. Number of observed topological states $n_{top}(Re)$ represented by COT representations as a function of Reynolds number. Inset: a least-square fitting of the scaling behaviour around the critical Reynolds number which is here assumed to be $Re_{c}=15400$.
4.2. Period estimation
One of the physical quantities that characterizes periodic motion is its period. The most common approach to determining the period is to apply a Fast Fourier Transform (FFT) to a time series of a physical observable such as kinetic energy and to analyse the resulting power spectrum. The period is estimated from the peak frequency of the power spectral density, as we did in Section 2.1 for $Re=14000$. In addition to such conventional physical observables, we here demonstrate that the period can also be estimated from time series of COT representations. More precisely, by computing the COT representation for each snapshot and arranging these descriptors in temporal order, we obtain a topology-based time series. Accordingly, we estimate characteristic periods from two distinct types of time series: (i) a continuous physical observable, the kinetic energy, and (ii) discrete signals derived from TFDA, given by the temporal evolution of COT descriptors. This framework enables a direct comparison between physics-based and topology-based descriptions of the dynamics. In particular, we will show that the estimated periods from the energy and the COT representation are quantitatively consistent, implying that the COT representation indeed reflects the essential physical information of the flow.
In addition to the spectral method (FFT), we employ two complementary period estimators. The first is based on the autocorrelation function (ACF); see Proakis (2007, Chapter 2). Autocorrelation measures the similarity between a signal and a time-shifted copy of itself as a function of the time lag; periodic signals produce peaks at lags corresponding to integer multiples of the period. For continuous-valued signals, we first rescale the data so that they have zero mean and unit variance (standard score, i.e., Z-score normalisation). For discrete COT-based signals, the symbolic descriptors are mapped to numerical values via a hash-based projection so that correlation operations can be performed. [^1] Because this projection depends on the specific hash mapping, the resulting correlations can vary; we retain this result here for reference. Dominant periods are then identified by detecting prominent peaks in the autocorrelation function.
The second estimator is based on Dynamic Time Warping (DTW) proposed by Sakoe and Chiba (1978). DTW evaluates the similarity between two temporal segments, while allowing for local stretching or compression along the time direction. To estimate candidate periods, the time series is divided into adjacent segments of a given length, each regarded as a cycle. For each candidate period, we compute the DTW distance between neighbouring segments in a constrained warping window (2 seconds). The mean warping distance over multiple shifted segmentations is then calculated, providing a quantitative measure of how consistently the temporal pattern repeats. A true period is expected to yield a small average warping distance, reflecting strong cycle-to-cycle similarity.
Table 3 summarises the estimated dominant periods for the lid-driven cavity flow, obtained from both kinetic energy and COT-based time series using the FFT, ACF and DTW estimators. Up to $Re=15250$, the estimated periods are consistent between observables and estimation methods. In this regime, the COT representation can be effectively regarded as a label identifying distinct phases within a single period. However, for $Re\geq 15300$, substantial discrepancies emerge between estimators and between signal types, indicating that the notion of a dominant period is no longer meaningful in this regime. A possible physical interpretation of this discrepancy is that the flow at $Re=15300$ is not strictly periodic. By applying the two estimation methods to two observables, the energy and the COT representation, we infer that the quasi-periodic regime observed at $Re=15500$ extends down to $Re=15300$. Consequently, the transition from the periodic state to the quasi-periodic state is likely to occur within the interval $15250<Re<15300$.
Table 3. Estimated dominant periods. For each Reynolds number, we report the period obtained from the kinetic energy time series and from a COT-based time series, using the ACF and DTW estimators. The periods estimated with the power spectral density of the energy, denoted by FFT, are listed up to $Re=15000$.
ReRe Energy (FFT) Energy (ACF) Energy (DTW) COT (ACF) COT (DTW) 14000 5.15 5.132 5.090 5.142 5.090 14250 5.13 5.147 5.137 5.147 5.076 14500 5.13 5.159 5.141 5.169 5.110 14750 6.45 6.360 6.373 6.370 6.191 15000 6.45 6.385 6.406 6.385 6.193 15250 6.397 6.426 6.386 6.244 15300 19.207 6.460 19.187 6.491 15350 19.298 6.761 19.156 6.367 15400 19.579 24.812 19.398 24.812 15450 19.718 24.758 19.406 12.907 15500 19.843 24.847 19.401 12.893 15550 19.872 24.753 19.449 24.753 15600 20.002 11.045 19.438 13.290 15650 20.013 10.634 19.560 13.195 15700 10.412 10.930 19.533 24.857 15750 20.220 10.108 19.476 24.773 15800 20.184 10.380 19.490 12.552 15850 10.348 10.880 19.520 24.743 15900 10.289 10.452 19.614 24.729 15950 20.342 9.983 19.607 24.837 16000 10.323 10.409 19.779 24.916
4.3. Energy and enstrophy variations with topological states
We consider the evolution of the kinetic energy $E(t)$ defined by (5) and that of the enstrophy defined by
(10) $Q(t)=\int_{\Omega}\frac{1}{2}(\omega(x,y,t))^{2}~{\mathrm{d}}x{\mathrm{d}}y.$
We here relate the topological states expressed by COT representations to the evolution of kinetic energy and enstrophy to identify their roles. In particular, we segment their time series according to topological states, which provides a new physical perspective of flow evolutions.
4.3.1. Periodic dynamics for $Re=14000$
Figure 14 shows the temporal variations of energy and enstrophy for $Re=14000$, which are periodic with the period $\tau_{p}=5.15$. They are segmented with colours according to the six topological states. We assign ID numbers to the topological states with the same COT representation in order of appearance. These IDs are the same as those introduced in Section 4.1. As shown in Figure 14(a), while energy increases, the topological state is mainly ID $1$, which is topologically equivalent to the flow pattern in Figure 8(a). Changes among the five remaining topological states occur while the energy decreases. This implies that the five topological states with IDs 2266 shown in Figure 8(b)–(f) are associated with relatively high energy dissipation. Indeed, enstrophy of these states is higher than average, as indicated in Figure 14(b). Notice that enstrophy is proportional to the energy dissipation rate in the bulk of the lid-driven cavity flow. From this behaviour of the simple flow we infer that the states with relatively many closed streamline structures in the corners have larger enstrophy or larger energy dissipation. This is in line with the physical intuition that the breakup of the corner vortices into smaller ones induces energy dissipation.
Figure 14. Variations of the flow for $Re=14000$; (a) energy and (b) enstrophy. They are segmented with the six COT representations for $0\leq t-T_{s}\leq 2.14\times\tau_{p}$, where the period is $\tau_{p}=5.15$. The same colour on the graph means the flow patten belongs to a set of topologically equivalent class with the same COT representation. The squares with labels aaff on the evolution represents the representative topological states shown in Figure 8(a)–(f). The topological state with ID $3$ appears for a short time. Note that some data are missing between the ID $5$ and 66 (around $t-T_{s}=4.2$) due to an error of the COT computation with psiclone.
4.3.2. Periodic dynamics for $Re=15000$
Figure 15(a) shows the energy variation which is color coded according to the topological states. The flow still exhibits a periodic behaviour in time with period $\tau_{p}=6.4$, but shows a qualitative change in the sequence of COT representations from the previous case. We observe nine topological states with IDs 1,2,5,8131,2,5,8-13 whose COT representations are listed in Table 4. The topological states with ID $1$, 22 and $5$ have been present for $Re=14000$. The qualitative change from the previous Reynolds number is that the state of ID $1$ is observed in the energy-increasing and energy-decreasing phases. This splitting behaviour appears as the transition diagram with the separated cycles in Figure 12(a). The central node connecting those cycles actually corresponds to the topological state with ID $1$.
Figure 15. Variations of the flow for $Re=15000$; (a) energy and (b) enstrophy for $0\leq t-T_{s}\leq 2.2\times\tau_{p}$, where the estimated period is $\tau_{p}=6.4$. The topological states with ID $1$, 22, 55 are the same as those appeared for $Re=14000$, while those with ID 881111 are observed for the first time. Data around $t-T_{s}=6,12$ is missing due to the error of the calculation of the COT.
According to the transition diagram, the right cycle $1\rightarrow 2\rightarrow 8\rightarrow 9\rightarrow 1$ corresponds to the energy-decreasing phase. This route corresponds to the local change in the flow topology in the top left corner, since we observe the change in the $c_{+}^{0}$ sequences of the COT representations in Table 4. Hence, the topological change in the eddy structures in the top left corner serves to reduce the energy. In contrast, Figure 15(a) also shows that the energy-increasing phase takes the route $1\rightarrow 5\rightarrow 10\rightarrow 11\rightarrow 10\rightarrow 12\rightarrow 13\rightarrow 1$, which appears as the left cycles of the transition diagram in Figure 12(a). We simply see the change in the sequence of $c_{+}^{1}$ for the topological states with ID $5$, $12$ and $13$ in Table 4. However, the COT representations for the topological states with ID $10$ and $11$ look different; therefore, we examine these flow patterns closely. Figure 16(a) is the streamline pattern at $t=t_{578}$ for $Re=15000$, which is identified as the topological state with ID $10$. Its COT representation contains the COT symbol $b_{--}\{\sigma_{-},\sigma_{-}\}$, indicating the existence of a figure-eight structure. This structure appears as a small self-connected saddle connection between the two boundary saddle connections of $c_{+}^{1}$ in the bottom left corner, as schematically shown in Figure 10(c). The streamline pattern at $t=t_{586}$ in Figure 16(b) is in the topological state with ID $11$. It has a small structure of nested saddle connections, which is represented by $b_{-+}(\sigma_{-},b_{-+}(\sigma_{-},\sigma_{+}))$ in the COT representation, between the two boundary saddle connections of $c_{+}^{1}$ in the bottom left corner. Hence, we conclude that the change in the eddy structures in the bottom left corner gives rise to the energy-increasing phase.
Figure 16. Streamline patterns at $t=t_{578}$ and $t=_{586}$ in the lid-driven cavity flow for $Re=15000$. They are identified as topological states with (a) ID $10$ and (b) ID $11$ respectively.
Another notable feature of Figure 15 is that the duration of the topological state of ID $5$ becomes longer than in the previous $Re=14000$ case. This is shown as a large node in the left cycle of the transition diagram in Figure 12(a).
Table 4. COT representations and IDs used in Section 4.1.
ID First appearance COT representation 1 $Re=14000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 2 $Re=14000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 5 $Re=14000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 8 $Re=15000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 9 $Re=15000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(b_{++}\{\sigma_{+},\sigma_{+}\})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 10 $Re=15000$ $\beta_{\emptyset-}(b_{--}\{\sigma_{-},\sigma_{-}\},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 11 $Re=15000$ $\beta_{\emptyset-}(b_{-+}(\sigma_{-},b_{+-}(\sigma_{+},\sigma_{-})),\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+})\})$ 12 $Re=15000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\},c_{-}(\sigma_{-}))\cdot c_{+}^{2}(\sigma_{+})\})$ 13 $Re=15000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+},c_{-}(\sigma_{-}))\cdot c_{+}^{2}(\sigma_{+})\})$ 33 $Re=15500$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(b_{++}\{\sigma_{+},\sigma_{+}\})\})$ 58 $Re=15500$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+},c_{-}(\sigma_{-}))\})$ 60 $Re=16000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(b_{++}\{\sigma_{+},\sigma_{+}\})\})$ 61 $Re=16000$ $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(b_{++}\{\sigma_{+},\sigma_{+}\})\cdot c_{+}^{2}(b_{++}\{\sigma_{+},\sigma_{+}\})\})$
4.3.3. Quasi-periodic dynamics for $Re=15500$
For $Re=15500$, the power spectrum of the energy time series indicates that the flow is considered quasi-periodic, as shown in Figure 3(b). We show the temporal variations of energy and enstrophy in Figure 17. We then observe thirty different topological states and complex transitions between them according to the transition diagram in Figure 12(b). However, in Figure 17(a), the route $1\rightarrow 2\rightarrow 9$ of the topological states, which has been observed for $Re=15000$, frequently appears around the local energy maxima. Also, around the local energy minima, we observe the topological state with ID $5$ that is the same as in the periodic regimes.
Since we have many topological states in the quasi-periodic regime, we focus on the five states that are most frequently observed in the time series, which are visualised as large nodes of the transition diagram in Figure 12(b). Four of them are the topological states that have appeared in periodic regimes; the first place is ID 55 ($0.30$), the second place is ID 11 ($0.24$), the fourth place is ID 99 ($0.091$), and the fifth place is ID 22 ($0.036$). The numbers in parentheses are the rate of occurrence, which will be discussed in detail later. The third most frequently observed topological state (ID $33$), which newly appears in this quasi-periodic regime, has the rate of occurrence $0.15$. Its COT representation is given by $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(b_{++}\{\sigma_{+},\sigma_{+}\})\})$ as shown in Table 4. This indicates that the change of the local topological structure in the bottom right corner, which has not been observed in periodic regimes, contributes to the increase in topological states. This topological state appears not only around the local maxima of enstrophy, but also around a middle value of enstrophy, leading to a splitting behaviour as shown in Figure 17(b).
Another newly observed topological state (ID $58$) with a relatively longer duration $0.012$ in this quasi-periodic regime has the COT representation $\beta_{\emptyset-}(\sigma_{-},\{c_{+}^{0}(\sigma_{+})\cdot c_{+}^{1}(\sigma_{+})\cdot c_{+}^{2}(\sigma_{+},c_{-}(\sigma_{-}))\})$. Since it appears mainly as minimum enstrophy, it can be compared with a similar enstrophy-minimum topological state (ID $10$) for $Re=15000$, which remains observed for $Re=15500$, but its occurrence rate decreases from $0.0306$ ($Re=15000$) to $0.0037$ ($Re=15500$). The difference in the COT representations indicates that the local streamline structures in the corners at the bottom are different; the local eddy structure is present in the bottom right corner for this topological state, while it is in the bottom left corner for the ID $10$ state. We infer that facilitation of eddy structures in the bottom right corner ($c_{+}^{2}$) is preferred in the enstrophy minima over having more structures in the bottom left corner for $Re=15500$.
Figure 17. Same as Figure 14 but for $Re=15500$ in $0<t-T_{s}\leq 110$. The COT representations for the topological states are listed in Table 4. (a) Variation of the energy. We observe the route $1\rightarrow 2\rightarrow 9$ around the energy maxima. (b) Variation of the enstrophy. The topological state with ID $33$ that is newly observed for $Re=15500$ appears several place with the high rate of occurrence. The topological state with ID $58$ appears around the local enstrophy minima with a relatively high rate of occurrence. Some data are missing due to the error of the calculation of the COT
4.3.4. Chaotic dynamics for $Re=16000$
[…]